### v2

# Beyond Prompting: Time2Lang – Bridging Time-Series Foundation Models and Large Language Models for Health Sensing

Abstract
Large language models (LLMs) show promise for health applications when combined with behavioral sensing data. Traditional approaches convert sensor data into text prompts, but this process is prone to errors, computationally expensive, and requires domain expertise. These challenges are particularly acute when processing extended time series data. While time series foundation models (TFMs) have recently emerged as powerful tools for learning representations from temporal data, bridging TFMs and LLMs remains challenging. Here, we present Time2Lang, a framework that directly maps TFM outputs to LLM representations without intermediate text conversion. Our approach first trains on synthetic data using periodicity prediction as a pretext task, followed by evaluation on mental health classification tasks. We validate Time2Lang on two longitudinal wearable and mobile sensing datasets: daily depression prediction using step count data (17,251 days from 256 participants) and flourishing classification based on conversation duration (46 participants over 10 weeks). Time2Lang maintains near constant inference times regardless of input length, unlike traditional prompting methods. The generated embeddings preserve essential time-series characteristics such as auto-correlation. Our results demonstrate that TFMs and LLMs can be effectively integrated while minimizing information loss and enabling performance transfer across these distinct modeling paradigms. To our knowledge, we are the first to integrate a TFM and an LLM for health, thus establishing a foundation for future research combining general-purpose large models for complex healthcare tasks.
Data and Code Availability
In this work, we use the StudentLife dataset (Wang et al., 2014) which is publicly available. Another dataset, the Major Depressive Disorder (MDD) dataset, is currently private, with plans for public release in the future. Code pertaining to public data can be made available upon request.
Institutional Review Board (IRB)
The study protocol for the major depressive disorder dataset has received full approval from the respective institutional review boards (IRB).
1 Introduction
Personal sensing devices such as mobile phones and wearables collect diverse behavioral data that can be used to assess mental health. Longitudinal data from these devices, including step counts and conversation patterns, effectively detect mental health conditions like depression (Nepal et al., 2024; Price et al., 2023). Meanwhile, large language models (LLMs) have recently demonstrated their ability to process a variety of data modalities beyond text, including audio (Ghosh et al., 2024) and images (Radford et al., 2021). A notable strength of LLMs is their in-context learning capability, enabling them to generalize from prompt-based examples without additional training (Dong et al., 2022). Building on these advancements, integrating behavioral sensing data with LLMs offers significant promise for developing personalized health models, with considerable impact for real-world applications (Cosentino et al., 2024).
Recent studies have incorporated health sensing data with LLMs to detect depression (Kim et al., 2024), atrial fibrillation (Liu et al., 2023b), and sleep quality (Cosentino et al., 2024). A common approach involves converting raw sensor data into text prompts to enable LLMs to process numeric time series data for classifying these conditions (Kim et al., 2024; Liu et al., 2023b). However, this method ignores the modality gap between sensing and text data (Spathis and Kawsar, 2024). Applying this technique to long-duration time series leads to higher error rates during prompting and requires a large amount of tokens (Section 2.4). Constructing effective prompts also requires substantial domain expertise, making it a challenging task even for domain experts (Section 4.2). Recent advancements in time series foundation models (TFM), such as Chronos (Ansari et al., 2024) and Moment (Goswami et al., 2024), have demonstrated their effectiveness in generating meaningful representations by capturing fundamental time series characteristics.
Therefore, we explore if incorporating a TFM in an LLM pipeline can address the aforementioned limitations. One way to efficiently incorporate two large models from different domains is through the concept of model re-programming (Chen, 2024). Previous works have successfully reprogrammed models from distinct domains such as text to biochemical sequence (Vinod et al., 2023) and speech to time series (Yang et al., 2021). Hence, we argue that incorporating a TFM into an LLM model through reprogramming can produce meaningful representations. Specifically, rather than using raw signals in prompts, we leverage TFMs as feature extractors (Figure 1). A key question is whether an LLM can understand TFM-derived features. To this end, we introduce an adapter that maps TFM features to LLM representations, effectively serving as a translator between the two models (Figure 1). We demonstrate our approach through behavioral sensing applications in mental health. Since behavioral sensing data is collected longitudinally in real-world settings, its unobtrusive nature results in long-duration sequences with realistic noise. Moreover, mental health is a high-stakes domain where minimizing errors, enhancing time-series comprehension, and improving efficiency are critical. Rather than proposing a new model, our goal is to maximize the potential of existing powerful models through an elegant adapter strategy. Thus, we explore whether two large models from distinct domains can be integrated with minimal information loss.
Toward the vision of seamless integration of time series data into LLMs, our contributions are as follows: (1) Time2Lang Framework. We propose Time2Lang, a framework that learns a mapping between two frozen foundation models (FMs): Chronos (Ansari et al., 2024) and LLaMA (Dubey et al., 2024) (Figures 1 and 3). To our knowledge, we are the first to integrate a TFM and an LLM for health tasks. Unlike most prior work, Time2Lang is trained within a self-supervised framework exclusively using intricate synthetic data generated from a Gaussian process (Section 3.1). (2) Evaluation on Longitudinal Mental Health Data. We assess Time2Lang’s performance on two real-world longitudinal mental health datasets with long-duration sequences. Specifically, we predict daily depression levels using step count data (17,251 days/labels) and analyze student flourishing from conversation duration (46 participants over 10 weeks) as described in Section 4. (3) Efficiency and Temporal Dynamics. We examine Time2Lang’s inference time, the impact of pre-training data size, the role of periodicity, and its relationship to classical time-series characteristics (Section 5). Our key findings (Section 6.1) are: (1) Time2Lang’s learned mapping minimizes information loss and enables positive transfer between two distinct domains—time series and text, (2) it maintains near constant inference time regardless of input length, and (3) its mapped LLM outputs align with fundamental time-series metrics, such as the autocorrelation factor. Overall, Time2Lang demonstrates that reprogramming can effectively integrate a TFM with an LLM, enhancing performance, efficiency, and time-series comprehension.
Figure 1: Time2Lang vs traditional prompting. An example of using sensing data to predict depression, comparing text-based prompting (left) with Time2Lang (right). In text-based prompting, the sensor signals are converted into text for LLM prompting. As an alternative, we introduce Time2Lang—our method learns a mapping ($f$ and $g$) between a TFM and an LLM, enabling health sensing tasks without the need for text conversion, while making the most of powerful models.
2 Related Work
2.1 Model Reprogramming
Model reprogramming is a training paradigm that efficiently repurposes a pre-trained, frozen large model across domains using an input transformation layer and an output mapping layer (Chen, 2024). Yang et al. (2021) reprogrammed a speech model for univariate time series classification, while Neekhara et al. (2022) proposed an adversarial approach to convert an image model for protein sequence prediction. Similarly, Vinod et al. (2023) focused on mapping a text model to biochemical sequences. From a reprogramming perspective, our work introduces several key advancements. It demonstrates that a TFM and an LLM can be effectively integrated for health-related tasks. Unlike most prior work, we adopt a self-supervised learning (SSL) framework instead of supervised learning. Additionally, our approach explicitly incorporates periodicity properties, further enhancing time-series understanding.
2.2 LLMs and Healthcare
Using LLMs to evaluate health outcomes has seen rapid growth (He et al., 2023; Han et al., 2023; Liu et al., 2025; Singhal et al., 2025). For instance, MedPaLM 2 (Singhal et al., 2025), an LLM fine-tuned for medical question answering has demonstrated strong performance by passing medical exam benchmarks. Similarly, models such as MedAlpaca (Han et al., 2023) and Me-LLaMA (Xie et al., 2024) also fine-tune general-purpose LLMs for comprehensive analysis of medical texts. For mental health tasks, Xu et al. (2024b) comprehensively evaluated different LLMs using online text data. Their work demonstrated that incorporating additional health context information in prompts enhances performance, but specialized domain knowledge is required.
2.3 LLMs and Sensor Data
Health sensing data consists of temporal numeric values that change over time. Approaches for time-series forecasting with LLMs involved converting the raw data into text prompts, as demonstrated by methods like LLMTime (Gruver et al., 2024) and PromptCast (Xue and Salim, 2023). For instance, Gruver et al. (2024) highlighted the importance of careful pre-processing, such as handling periods (“.”) and spaces (“ ”), to ensure effective tokenization of floating-point numbers. Recently, alternative approaches have used multi-modal LLMs that encode sensing data as images Yoon et al. (2024) or use modality-specific encoders Belyaeva et al. (2023). For example, Yoon et al. (2024) proposed transforming time-series data into visual prompts for use with vision-language models.
2.4 Limitations of integrating mobile sensing with LLMs
Figure 3: Time2Lang Framework. To meaningfully integrate Timeseries Foundation Models (here: Chronos $C$) and Large Language Models (LLaMA $M$), we train two smaller networks $f$ and $g$ that optimally map TFM features ($\mathbf{z^{c}}$) to an LLM. The learned embeddings from $f$ and $g$ are $\mathbf{z^{i}}$ and $\mathbf{z^{o}}$, respectively. To improve positive knowledge transfer, we use a residual connection between the TFM and LLM features ($\mathbf{z^{c}}\rightarrow\mathbf{z^{m}}$) only during training.
Despite advances for integrating sensor data into LLMs, the most common approach remains incorporating the raw sensor data as text strings into prompts without extensive pre-processing (Kim et al., 2024; Liu et al., 2023b; Xu et al., 2024a). While this technique has shown promise for aggregated short-duration time series (e.g., $<$ 200 time steps), behavioral sensing data often has long duration sequences. For example, one day of data in one minute intervals consists of 1440 data points. Sampling rates $>1$Hz will have substantially more data points. Therefore, representing sensor data as text has the following limitations. First, an LLM performance decreases with longer duration sequences (Yoon et al., 2024). To empirically demonstrate this using LLaMA 3.2 (Dubey et al., 2024), we perform a mean prediction task using synthetic data, where LLaMA is asked to predict the mean with increasing sequence length. In Figure LABEL:fig:limitations, we notice that the mean absolute error (MAE) increases with increasing sequence length. Second, long duration sequences require significantly more tokens, which may quickly exhaust the models limits. Specifically, observe that the number of tokens required for LLaMA 3.2 is 10$\sim$ 10$\times$ 10 × the length of the time series (Figure LABEL:fig:limitations). Third, effective prompting is challenging and requires strong domain knowledge (adding specific knowledge improves prompting performance – Table LABEL:tab:overall_performance LLaMA-ICL vs LLaMA-ICL-DK), remaining an active research area in NLP (Zamfirescu-Pereira et al., 2023; Bommasani et al., 2021; Liu et al., 2023a).
Our approach offers an alternative perspective for addressing these limitations with mental health as an application. By learning a mapping that seamlessly integrates a TFM into the LLM workflow, Time2Lang inputs fixed-size embeddings directly into the LLM, eliminating the need for prompting. We show that the Time2Lang framework is more scalable and efficient compared to generative LLM prompting, while preserving the performance of the TFM.
3 Methods
3.1 Synthetic Data Generation
Human behavioral signals often display periodic patterns, and leveraging self-supervised learning (SSL) to capture these patterns has been shown to produce robust representations (Yang et al., 2022; Gu et al., 2019). To enable Time2Lang to learn periodicity, we generate synthetic data that mimics intricate sensing patterns by sampling from a Gaussian Process (GP) regression. In particular, we modify KernelSynth (Ansari et al., 2024) to produce time series that exactly repeats itself after specific interval. A GP is defined by a mean function $m(a)$ (set to 0) and a covariance function or kernel $k(a,a^{\prime})$ which models the joint variability of the pair of points $(a,a^{\prime})$ (Schulz et al., 2018).
We perform the following steps to generate the synthetic data. First, we randomly sample up to $j=4$ non-periodic kernels from a bank $\mathcal{K}$ consisting of the following kernels: constant, white noise, linear, radial basis function, and rational quadratic ($\{k_{1}(a,a^{\prime}),\cdots,k_{j}(a,a^{\prime})\}$). These kernels are useful to model general sensor patterns such as trend, noise, and local variation. Second, we sample a periodic value $p$ from a set of pre-defined values $p\in\{30,60,90,120,150,180\}$, and generate the exponential sine squared kernel, $k_{per}(a,a^{\prime})$. This infuses seasonality into the generated data such that patterns exactly repeat themselves. Third, given a set of non-periodic and periodic kernels $\{k_{1}(a,a^{\prime}),\cdots,k_{j}(a,a^{\prime}),k_{per}(a,a^{\prime})\}$, we compose them into a single $k(a,a^{\prime})$ by applying a random binary operation $+$ or $\times$. Finally, a synthetic time series of length $l=1440$ (one day of data in one-minute intervals) is generated by sampling from the GP with $m(a)=0$ and covariance function $k(a,a^{\prime})$.
3.2 Time2Lang
Time2Lang consists of four sequential components (Figure 3), a frozen TFM ($C$), a learnable input encoder ($f$), a frozen LLM ($M$), and a learnable projection ($g$). It is worth noting that inputs to the input encoder and the projection are the outputs of $C$ and $M$, respectively.
3.2.1 Training
We train Time2Lang within an SSL framework using synthetic data, where periodicity prediction is the pretext task. Formally, given a time series $\mathbf{x}\in\mathbb{R}^{n}$ of length $n$ and the corresponding periodicity $y\in\{30,60,90,120,150,180\}$, the objective is to learn $f$ and $g$ to accurately predict periodicity as a multi-class classification problem.
Time2Lang is trained following the steps outlined in Algorithm 1. Given an pair $(\mathbf{x},y)$, we first extract features from $\mathbf{x}$ using a frozen TFM - Chronos ($C$) (Ansari et al., 2024). The resulting Chronos embeddings $\mathbf{z^{c}}\in\mathbb{R}^{c\times d}$ are temporal features that are fixed irrespective of input size, where context length $c=513$ and features $d=768$. Previous work has shown that TFMs, such as Chronos, effectively extract meaningful time-series features. Instead of using raw signals as prompts, we leverage these extracted features. However, since there is no guarantee that an LLM can interpret these features, we aim to explicitly establish a meaningful mapping between TFM-derived features and an LLM, specifically LLaMA 3.2 ($M$). To achieve this, we train an input encoder ($f$), a lightweight 1D ResNet-style CNN. This encoder processes the transposed output of $C$, $\mathbf{z^{c}}^{T}$, generating compressed temporal features $\mathbf{z^{i}}\in{\mathbb{R}^{d\times c}}$, where the feature dimension $d$ and sequence length $c$ are determined by the convolution and pooling parameters ($d=65$, $c=64$ in our case). The encoded representation, $\mathbf{z^{i}}^{T}$, is passed to input embedding of LLaMA ($M$), instead of token ids. The output is averaged to obtain $\mathbf{z^{m}}\in\mathbb{R}^{2048}$. Subsequently, we project $\mathbf{z^{m}}$ using two fully-connected (FC) layers with 768 and 256 neurons. Additionally, a residual connection is introduced between the averaged $\mathbf{z^{c}}$ and the first FC layer to enhance information retention from Chronos. Finally, $\mathbf{z^{o}}$ is passed through a output FC layer $l$ to produce a logit vector $\mathbf{\hat{y}}$. Conceptually, the input encoder ($f$) and projection network ($g$) are trained to optimize the probing of the frozen LLaMA model, leveraging periodicity as a pretext task to learn useful embeddings. The model is trained end-to-end by optimizing the multi-class cross entropy loss as follows:
$\mathcal{L}=-\frac{1}{N}\sum_{k=1}^{N}\log\frac{\exp{(\mathbf{\hat{y}}_{k,y_{k}})}}{\sum_{j=1}^{6}\exp{(\mathbf{\hat{y}}_{k,j})}}$ (1)
where $N$ is batch size, $y_{i}$ is the correct index for sample $i$, and $6$ represents the number of periodic classes.
Input: $N$ synthetic time series with periodicity $\{\mathbf{x}_{k},y_{k}\}_{k=1}^{N}$, TFM: $C$, LLM: $M$, Input Encoder: $f$, Projection: $g$, Output layer: $l$.
for $k\in\{1,\dots,N\}$ do
       $\mathbf{z^{c}}_{k}\leftarrow C(\mathbf{x}_{k})$
       $\mathbf{z^{i}}_{k}\leftarrow f((\mathbf{z^{c}}_{k})^{T})$
       $\mathbf{z^{m}}_{k}\leftarrow M((\mathbf{z^{i}}_{k})^{T})$
       $\mathbf{z^{m}}_{k}\leftarrow\text{mean}(\mathbf{z^{m}}_{k})$
       $\mathbf{z^{o}}_{k}\leftarrow g(\mathbf{z^{m}}_{k},\text{mean($\mathbf{z^{c}}_{k}$))}$) # residual
       $\mathbf{\hat{y}}_{k}\leftarrow l(\mathbf{z^{o}}_{k})$
end for
Compute $\mathcal{L}=-\frac{1}{N}\sum_{k=1}^{N}\log\frac{\exp{(\mathbf{\hat{y}}_{k,y_{k}})}}{\sum_{j=1}^{6}\exp{(\mathbf{\hat{y}}_{k,j})}}$
Update networks $f$ and $g$ to minimize $\mathcal{L}$.
Algorithm 1 Time2Lang Training
Our rationale for choosing Chronos and LLaMA is based on several key considerations. First, Chronos draws inspiration from language modeling, specifically, its approach to time series quantization and discretization is analogous to tokenization in language models. Second, it transforms variable-length inputs into fixed-size embeddings, enhancing efficiency as input lengths grow. Lastly, these models are well-established within their respective fields, promoting the development of findings that are more broadly generalizable.
3.2.2 Implementation
We generate 200K synthetic samples, as outlined in Section 3.1. These samples are further split into training, validation, and testing set in 70/10/20 ratio. Time2Lang is trained exclusively on synthetic samples for 25 epochs, using a batch size of 16, on an NVIDIA A4500 GPU. The training process optimizes the model to predict periodicity by minimizing the cross-entropy loss (Equation 1). Specifically, the Adam optimizer with a learning rate of $5\times 10^{-4}$ is used to update the parameters of the networks $f$ and $g$. Given the sensitive nature of mental health data, we implement the model locally, utilizing Chronos base and LLaMA 3.2 as frozen foundation models (FMs) with 200M and 1B parameters, respectively. Moreover, if our approach works for these smaller models, it will improve with larger models due to scaling laws (Ansari et al., 2024; Kaplan et al., 2020). The input encoder $f$ consist of 300K parameters ($\approx 0.03\%$ of LLaMA), while the projection $g$ consists of 1.7M parameters ($\approx 0.17\%$ of LLaMA).
4 Experimental Setup
In this section, we describe our rationale for choosing datasets, baselines, and evaluation protocol.
4.1 Datasets
We sought to evaluate Time2Lang on diverse longitudinal datasets in terms of sensing modality, devices, population, and tasks. Therefore, we examine: (1) daily depression detection in a clinical population using step count data from a wearable device, and (2) classification of flourishing levels in a student population (non-clinical) based on their conversation duration from mobile phones over the course of an academic term. We describe dataset information relevant to our analyses below.
4.1.1 Major Depressive Disorder (MDD)
The Major Depressive Disorder (MDD) longitudinal study investigates intra-day fluctuations in depressive symptoms among individuals diagnosed with MDD. In this 90-day study, we recruited participants aged 18 years and older, residing in the United States. To determine eligibility, we administered the structured clinical interview for DSM-5 (SCID), ensuring that only individuals with MDD, but without co-morbid bipolar disorder, active suicidality, or psychosis, were enrolled. Each participant was instrumented with a Garmin vivoactive 3 wearable and installed our Android application. While the Garmin passively collected sensing data, the mobile application prompted to the user to answer the Patient Health Questionnaire-9 (PHQ-9) three times a day (Kroenke et al., 2001). This self-reported survey consisting of nine questions rated on a 0-3 Likert scale is well-validated tool to evaluate depression. To encourage engagement, participants received \$1 per completed survey, with an additional \$50 bonus for maintaining a survey completion rate greater than 90%. Now, we describe information pertaining to the analysis in this work.
Physical activity plays a crucial role in depression fluctuations, and step count provides a straightforward measure of activity levels. Notably, a daily step count exceeding 5,000 has been associated with fewer depressive symptoms (Bizzozero-Peroni et al., 2024). Hence, we assess daily depression levels using step count data from the Garmin wearable. It is worth noting that daily depression prediction is challenging problem compared to longer time-scales. In our analysis, the input sensing data consists of 1,440 minute-level step count values per day, while the ground truth is derived from the Patient Health Questionnaire-9 (PHQ-9), where a score of 10 or higher indicates depression (Figure 5). To handle missing data, we apply a threshold-based approach: if missingness is below 25%, we impute missing values using zero-filling; otherwise, we exclude the day from the analysis. The PHQ-9 score is averaged within each day to align with the step count data. Our dataset comprises 256 participants, covering a total of 17,251 person-days (68% depressed days), with corresponding depression labels.
4.1.2 StudentLife
The StudentLife study (Wang et al., 2014) tracked the behavior of 48 undergraduate students at Dartmouth College over a 10-week term using mobile sensing. An Android application passively collected conversation duration from ambient sound, while self-reported surveys were administered to examine students’ mental health. Beyond academics, psychological well-being is an important part of college life, involving personal growth, social connections, and civic engagement. The Flourishing Scale (FS) (Diener et al., 2010) measures this well-being through an 8-item questionnaire rated on a 1-7 scale from strongly disagree to strongly agree (Appendix C). The total FS score provides a single measurement of psychological well-being. As social contact is directly examined in FS, we focus on classifying students into high and low flourishing (median split) groups based on their conversation duration throughout the academic term. Specifically, the input is a variable-length conversation duration time series up to 10 weeks, where each data point represents hourly conversation duration, and the output is a binary indicator of high or low flourishing. We use data from 46 students with and average time series length of $1455\pm 193$.
Table 1: Time2Lang performance comparison against baselines. Values in bold denote when models outperform the upper bound (Chronos). Higher values indicate better performance.
Model Depression (step count) Flourishing (conversation)
AUROC AUPRC AUROC AUPRC
Random (Lower) 0.50 0.68 0.50 0.52
Chronos (Upper) 0.56 (0.01) 0.72 (0.01) 0.70 (0.15) 0.69 (0.14)
[…]
LLaMA-ICL-DK 0.53 (0.03) 0.68 (0.02) 0.54 (0.10) 0.63 (0.15)
Chronos + LLaMA 0.54 (0.01) 0.70 (0.00) 0.56 (0.16) 0.65 (0.15)
Time2Lang 0.57 (0.01) 0.73 (0.00) 0.71 (0.10) 0.75 (0.11)
4.2 Baselines
We evaluate Time2Lang using several baseline approaches. The lower bound baseline is Random classification where AUROC is 0.5 and AUPRC is the proportion of positive samples. As we transfer information from Chronos to LLaMA, we treat Chronos as the upper bound baseline, assuming some information loss occurs during translation. To determine whether reprogramming through a mapping function is necessary, we directly embed Chronos into the LLaMA embedding without a mapping function (with embeddings padded as needed), referred to as Chronos + LLaMA. To compare against prior prompting methods, we use a 3-shot in-context learning (ICL) approach with LLaMA, referred to as LLaMA-ICL. The prompt design follows established best practices from existing literature. First, from LLMTime (Gruver et al., 2024), we round numerical values to two-digit precision, remove periods without adding spaces, and enclose the time series values in quotes to convert them into strings (e.g., $0.123\rightarrow``012"$). Second, building on HealthLLM (Kim et al., 2024), we include health-related context (e.g., “Some symptoms of depression include a lack of motivation and fatigue, which can be inferred from daily step counts.”) and temporal context (e.g., “{20, 25, …, 30”}). To further support LLaMA-ICL, we incorporate domain-specific knowledge – LLaMA-ICL-DK, which is crucial for guiding the model effectively. For example, “Conversation duration is associated with flourishing, with longer durations typically indicating higher levels of flourishing.” This addition from the self-reported instrument provides more context to better understand and interpret the data within the specific application domain.
4.3 Evaluation
After initial pre-training, Time2Lang can be used for a wide variety of different time-series tasks (here convo and steps) without further fine-tuning. For downstream evaluation, we process the input sensing data from real datasets and extract embeddings from the penultimate linear layer without the residual connection to determine whether the information from Chronos has been effectively distilled. The extracted embeddings are then used to train either a logistic regression (LR) or random forest (RF) model for mental health classification. Prior to training, we split the dataset into an 80/20 subject-wise split for training and hold-out testing. Next, we perform grid search cross-validation (3 folds) on the training data to identify the optimal model parameters. Once the best parameters are selected, we evaluate the final model on the hold-out test set using the area under the receiver operating characteristic curve (AUROC) and the area under the precision-recall curve (AUPRC). AUROC evaluates the overall discriminative power of the model, while AUPRC places greater emphasis on the performance for the positive class. Moreover, we repeat the experiments across five different shuffled training splits and report the standard deviations of the results for the best model (LR or RF).
5 Results
5.1 Overall Performance
Table LABEL:tab:overall_performance presents the performance comparison of Time2Lang with the different baselines for depression detection and flourishing classification. For depression detection using step count, Time2Lang achieves an AUROC of 0.57 and an AUPRC of 0.73, outperforming the baselines. Compared to Chronos, Time2Lang slightly improves the AUROC (0.56 vs. 0.57) and AUPRC (0.72 vs. 0.73) for depression detection. For flourishing classification using conversation duration, Time2Lang surpasses the prompting-based and Chronos + LLaMA baselines, with AUROC and AUPRC improvements ranging from 0.15 to 0.20 and 0.10 to 0.14, respectively. Compared to Chronos, Time2Lang achieves an AUROC of 0.71 (+0.01) and an AUPRC of 0.75 (+0.06). The higher AUPRC suggests that Time2Lang improves positive case detection. Overall, Time2Lang consistently outperforms most baselines, demonstrating marginal gains over Chronos in three cases and a more substantial improvement in one case.
Some other interesting observations from Table LABEL:tab:overall_performance are as follows. First, incorporating domain knowledge into prompting enhances performance across tasks in both AUROC and AUPRC (LLaMA-ICL vs. LLaMA-ICL-DK). Notably, for depression detection, LLaMA-ICL underperforms the lower-bound baseline, with an AUROC of 0.49 and an AUPRC of 0.67, suggesting that prompting for long-duration sequences poses challenges. However, adding domain knowledge improves AUROC by 0.04 and AUPRC by 0.01, demonstrating its positive impact. Second, integrating Chronos with LLaMA leads to a decline in performance, particularly in the flourishing prediction task, where AUROC and AUPRC drop by 0.14 and 0.04, respectively, compared to Chronos alone. In contrast, Time2Lang enables positive performance transfer, a 0.01 increase in AUROC while achieving a notable 0.06 increase in AUPRC, highlighting the necessity of having a mapping to bridge two distinct domains.
5.2 Efficiency Analysis
To evaluate the efficiency of Time2Lang and LLaMA prompting, we assess inference times using two approaches. First, we measure inference time across varying time series lengths: $\{256,512,1024,2048,4096\}$ (Figure LABEL:fig:efficient(a)). Second, given that the StudentLife conversation data represents a variable-length time series, we evaluate inference time across all data samples (Figure LABEL:fig:efficient(b)). Prior to measurement, we ensure consistent execution by transferring the data and model to the device, synchronizing GPU, and performing a warm-up phase with 100 iterations. As shown in Figure LABEL:fig:efficient, Time2Lang consistently achieves lower inference times compared to prompting across different data lengths in the synthetic experiment and conversation duration data in the StudentLife dataset. Notably, the inference time for Time2Lang remains relatively stable as the data length increases, whereas LLaMA prompting exhibits a proportional increase in inference time. This disparity is primarily due to the increasing relationship between input token length and sensing data length, which affects the computational overhead of prompting-based approaches.
5.3 Effect of Parameters
We evaluate the impact of synthetic data size and number of periodicity classes on downstream performance. Specifically, we consider dataset sizes of 50K, 100K, and 200K samples and periodicity class counts of 3 (30 to 90), 6 (30 to 180), and 9 (30 to 270), resulting in nine combinations. As shown in Figure LABEL:fig:effect_of_pre_training, pre-training with 200K samples and 6 periodicity classes achieved the best performance across all cases. Among dataset sizes, performance improves as data increases from 50K to 100K, but the difference between 100K and 200K is marginal. Similarly, increasing periodicity classes from 3 to 6 generally enhances performance. However, further increasing the classes to 9 does not yield additional improvements, regardless of dataset size. We attribute this plateau to the model’s parameter capacity, as a 9-class classification problem demands a more complex network, which may exceed the capacity of the trainable parameters.
5.4 Case Study: Correlation of Embeddings and ACF
We aimed to evaluate whether, after reprogramming through Time2Lang, the LLM outputs effectively capture fundamental time series characteristics. Thus, we computed the Spearman rank correlation between the Auto-Correlation Factor (ACF) of the original time series and the embeddings. First, we calculated the ACF with 10 lags from the original sensing data. Then, we computed the pairwise correlation between each ACF lag and the corresponding Time2Lang embedding across the samples. The heatmaps in Figure LABEL:fig:correlation (a) and (c) show that Time2Lang embeddings capture both positive and negative correlations across conversation and step count data. Notably, we observe stronger correlations at higher time lags (7–9), suggesting that Time2Lang embeddings retain information extending beyond recent past observations. To further analyze the strength of these correlations, we identified the highest correlation value for each embedding dimension at any lag. The distribution of these values, shown in Figure LABEL:fig:correlation (b) and (d), indicates that for conversation data, 174 dimensions ($\sim$68%) have at least one lag where the correlation exceeds 0.3, signifying a moderate correlation. Similarly, for step count data, 70 dimensions ($\sim$27%) demonstrate a moderate correlation greater than 0.3.
6 Discussion
6.1 Findings
Time2Lang achieved an AUROC=0.57 and AUPRC=0.73 for predicting daily depression based on step count and an AUROC=0.71 and AUPRC=0.75 for predicting flourishing. In the context of mental health, its depression prediction performance aligns with findings from previous studies (Xu et al., 2023). Among prompting approaches, incorporating specific domain knowledge enhances performance, consistent with prior observations by Kim et al. (2024). However, Time2Lang consistently outperforms prompting across both metrics and tasks. The LLaMA + Chronos approach showed markedly lower performance than Chronos alone, indicating that information is lost when integrating two distinct models. Our goal is to minimize performance loss during translation when leveraging Chronos. In this regard, Time2Lang outperforms LLaMA + Chronos, demonstrating that learning a direct mapping yields better results than simply combining embeddings. Ultimately, Time2Lang achieves both marginal and notable improvements over Chronos, showcasing that a TFM can be integrated with an LLM by minimizing information loss across these two distinct domains.
Through our efficiency analysis, we find that Time2Lang’s inference time is steady across increasing length at approximately 0.04 seconds. However, we notice that LLaMA Prompting increases with longer duration sequences primarily because of more tokens. The faster inference time of Time2Lang can be attributed to the TFM that produces a fixed sized embeddings, irrespective of input length. Our correlation analysis reveals that after mapping, the LLM outputs from Time2Lang exhibit moderate to high correlations with ACF, indicating that our method effectively captures key time series properties. We attribute this capability to our self-supervised pre-training, which is designed to predict periodicity. Interestingly, this pre-training on synthetic data, generalizes to real-world datasets, demonstrating the robustness of our approach.
6.2 Implications
Our work has significant implications for the development of personalized mental health LLMs and context-aware interventions. It represents an essential first step toward integrating sensing data into LLM-driven mental health systems, enabling them to provide critical, user-specific insights. Prior research has demonstrated that mental health symptoms fluctuate significantly within a single day and across different populations (Lim et al., 2018; Crowe et al., 2019). A personalized system that continuously monitors behavioral sensing data can identify individual-specific patterns, offering a more nuanced understanding of mental health variations. Unlike traditional check-ups, which provide only periodic assessments, continuous monitoring enables dynamic, real-time recommendations tailored to the user’s needs. However, to deliver accurate and effective mental health recommendations, the first step is ensuring that the model can comprehend and process sensing data efficiently. Scalability and computational efficiency are crucial for handling continuous sensing streams. As discussed in Sections 5 and 6.1, the Time2Lang approach minimizes information loss while achieving performance comparable to a TFM, improving efficiency, and establishing strong connections to fundamental time series characteristics. These results show that fusing two FMs from distinct domains can lead to meaningful representations, consequently paving the way for future work that may enable real-time generation and dynamic mental health interventions for users.
6.3 Limitations & Future Work
We acknowledge the limitations of our work, particularly its reliance on uni-modal time series data, as Time2Lang (Chronos) processes a single modality. Extending it to multi-modal data is a promising direction that would enhance Time2Lang’s ability to handle richer data. While we demonstrate Time2Lang’s effectiveness for mental health, its potential extends to other areas including physiological sensing, industrial sensors, and environmental monitoring. Therefore, we plan to evaluate our method on more diverse datasets to evaluate generalizability in the future.
The selection of Chronos and LLaMA is motivated by their fixed-length embeddings, relation to language modeling, and practical utility, as discussed in Section 3.2. However, different tasks or modalities may benefit from alternative pre-trained models. Rather than prescribing a one-size-fits-all approach, we aim to provide a new perspective on leveraging LLMs for sensing data analysis. Future work should systematically evaluate modality interactions and model suitability.
Designing computationally efficient reasoning methods for behavioral sensing in mental health remains challenging. Prior research suggests reasoning is strongly tied to scaling (Moon et al., 2024; Girdhar et al., 2023; Spathis and Kawsar, 2024). While our approach emphasizes efficiency, its reasoning capabilities remain limited. However, Time2Lang provides a foundation for integrating time series data into LLMs, which we aim to extend for more complex reasoning and generative tasks. Ultimately, our goal is to develop a personalized LLM that understands individual user behavior through daily data, enabling more tailored and insightful interactions.
7 Conclusion
In this work, we propose Time2Lang, a framework that integrates sensing data with LLMs by learning a direct mapping from a TFM (Chronos) to an LLM (LLaMA). This approach overcomes the limitations of text-based prompting. Through extensive evaluation on real-world longitudinal datasets for tasks such as daily depression prediction and flourishing classification, we demonstrate that Time2Lang positively integrates a TFM and an LLM while maintaining stable inference times across varying input lengths. Furthermore, our correlation analysis highlights Time2Lang effectively reprograms LLM outputs to capture fundamental time-series properties such as auto-correlation. Ultimately, Time2Lang enhances efficiency, accuracy, and time-series comprehension compared to traditional prompting methods. By bridging the gap between behavioral sensing and LLM-based inference, Time2Lang represents a significant step toward enhancing AI-driven mental health assessment. Future work will explore its extension to reasoning with multi-modal sensing data, further improving its applicability in personalized mental health monitoring and intervention systems.
Appendix A Architecture & Hyperparameters
The input encoder ($f$) comprises 13 convolutional layers (Table 2): an initial layer, followed by six 2-layer blocks (Tables 4 and 4). The input channels are 768, and we begin with 32 base filters, using a kernel size of 3 and stride of 2. The projection layer consists of two fully connected layers with 768 and 256 features, respectively. Each layer is followed by batch normalization and ReLU activation to ensure stable training and improved representation learning.
Table 2: ResNet-Style CNN
Layer
Conv1D
BatchNorm
ReLU
(Basic Block Type 1) x 1
(Basic Block Type 2) x 5
BatchNorm
ReLU
Table 3: Basic Block Type 1
Layer
Conv1D
BatchNorm
ReLU
Dropout
Conv1D
Table 4: Basic Block Type 2
Layer
BatchNorm
ReLU
Dropout
Conv1D
BatchNorm
ReLU
Dropout
Conv1D
Maxpool
In our downstream evaluation, logistic regression performed the best for depression detection using step count, whereas random forest is used for classifying flourishing using conversation duration. We obtained the best model using the following hyperparameters. Logistic regression: {’penalty’: [’l1’, ’l2’, ’elasticnet’, ’none’], ’C’: [0.1, 1, 10, 100], ’solver’: [’saga’], ’l1_ratio’: [0.5, 0.7, 0.9]} and Random Forest: {’n_estimators’: [50, 100], ’max_depth’: [10, 20, 50], min_samples_split’: [2, 10], ’min_samples_leaf’: [1, 4], ’max_features’: [’sqrt’, ’log2’, None], ’bootstrap’: [True, False]}
Appendix B Data
B.1 MDD
The MDD study recruited 300 participants for a 90-day study. The sample is predominantly middle-aged, with an average age of 40.13 years and an age range of 19 to 79 years, covering early adulthood to late older adulthood. The majority of participants are White (79%), heterosexual (66%), and women (84%), with an average age of 40 years. The racial distribution of minority groups, including Black, Asian, and American Indian or Alaska Native participants, closely reflects the U.S. MDD population, while Hispanic and Latino individuals (12%) match national demographic proportions. Most participants (93%) have some college experience, with 68% holding a college degree, and over half (61%) are employed. Household incomes are distributed across all income brackets, largely aligning with the general U.S. distribution at the time of sampling. The study also maintained high engagement, with an average compliance rate of 82.07%, based on daily Ecological Momentary Assessment (EMA) responses per participant over the 90-day period.
Figure 5: Label statistics. PHQ-9 Distribution in the MDD dataset. Scores $\geq$ 10 are considered depressed. 68% of the days correspond to depression.
Appendix C Self-reported Surveys and Ecological Momentary Assessments
C.1 Patient Health Questionnaire - 9 (PHQ-9) (Kroenke et al., 2001)
In our MDD study, the user is asked over the past four hours if they have experienced the questions/statements in Table 5.
Table 5: Patient Health Questionnaire - 9 (PHQ-9).
No. Statement
1 Little interest or pleasure in doing things
2 Feeling down, depressed, or hopeless
3 Trouble falling or staying asleep, or sleeping too much
4 Feeling tired or having little energy
5 Poor appetite or overeating
6 Feeling bad about yourself or that you are a failure or have let yourself or your family down
7 Trouble concentrating on things, such as reading the newspaper or watching television
8 Moving or speaking so slowly that other people could have noticed
9 Thoughts that you would be better off dead, or of hurting yourself
C.2 Flourishing Scale (Diener et al., 2010).
The user to indicate their agreement with each item in Table 6 by rating each statement from 1 to 7: (1) strongly disagree, (2) disagree, (3) slightly disagree, (4) mixed or neither agree nor disagree, (5) slightly agree, (6) agree, (7) strongly agree.
Table 6: Flourishing scale statements.
No. Statement
1 I lead a purposeful and meaningful life.
2 My social relationships are supportive and rewarding.
3 I am engaged and interested in my daily activities
4 I actively contribute to the happiness and well-being of others
5 I am competent and capable in the activities that are important to me
6 I am a good person and live a good life
7 I am optimistic about my future
8 People respect me
Appendix D Prompt Templates
D.1 Depression Prompt
The prompt template for depression is given below. In this template, we use temporal and health context. Furthermore, we convert the labels from 0 to 1 (as a probability). This approach produced better results than directly predicting 0 or 1.
You are an AI assistant that evaluates if a user is depressed based on their step count in a day. Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request in the format provided. Task: Some symptoms of depression are a lack of motivation and fatigue. It is linked to the number of steps taken in a day. Using the step count time series for a day, output a probability between 0 to 1 where 1 indicates high depression and 0 indicates low depression Input: - Step Count: {<<<”sensor data”>>>}
D.2 Depression Prompt with Knowledge
The prompt template for depression is given below where add domain knowledge (highlighted).
You are an AI assistant that evaluates if a user is depressed based on their step count in a day. Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request in the format provided. Task: Some symptoms of depression are a lack of motivation and fatigue. It is linked to the number of steps taken in a day such that higher step count is associated with lower depression symptoms. Using the step count time series for a day, output a probability between 0 to 1 where 1 indicates high depression and 0 indicates low depression Input: - Step Count: {<<<”sensor data”>>>}
D.3 Flourishing Prompt
The prompt template for flourishing is given below. In this template, we use temporal and health context.
You are an AI assistant that evaluates if a user is flourishing based on their conversation duration over time. Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request in the format provided. Task: We know that conversation duration is related to flourishing. Using the conversation duration (seconds) time series given below, output a probability between 0 to 1 where 1 indicates high flourishing and 0 indicates low flourishing. Input: - Conversation duration: {<<<”sensor data”>>>}
D.4 Flourishing Prompt with Domain Knowledge
The prompt template for flourishing is given below. In this template, we use temporal and health context with domain knowledge.
You are an AI assistant that evaluates if a user is flourishing based on their conversation duration over time. Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request in the format provided. Task: We know that conversation duration is related to flourishing such that higher conversation duration is linked to higher flourishing. Using the conversation duration (seconds) time series given below, output a probability between 0 to 1 where 1 indicates high flourishing and 0 indicates low flourishing. Input: - Conversation duration: {<<<”sensor data”>>>}